### v3

# Correlated equilibrium implementation: Navigating toward social optima with learning dynamics

Abstract.
Implementation theory has made significant advances in characterizing which social choice functions can be implemented in Nash equilibrium, but these results typically assume sophisticated strategic reasoning by agents. However, evidence exists to show that agents frequently cannot perform such reasoning. In this paper, we present a finite mechanism which fully implements Maskin-monotonic social choice functions as the outcome of the unique correlated equilibrium of the induced game. Due to the results in (Hart and Mas-Colell, 2000), this yields that even when agents use a simple adaptive heuristic like regret minimization rather than computing equilibrium strategies, the designer can expect to implement the SCF correctly. We demonstrate the mechanism’s effectiveness through simulations in a bilateral trade environment, where agents using regret matching converge to the desired outcomes despite having no knowledge of others’ preferences or the equilibrium structure. The mechanism does not use integer games or modulo games.
conference: ; ;
1. Introduction
Implementation theory has made significant strides in characterizing implementable social choice functions. (Maskin, 1999) establishes that any social choice function satisfying monotonicity can be implemented in Nash equilibrium. However, this and many subsequent results rely on agents being able to compute and play Nash equilibria (and indirectly, best responses) - an assumption that has faced increasing scrutiny from both theoretical and empirical perspectives (see for instance (Goeree and Louis, 2021), (Li, 2017)).
This has led to a growing interest in implementation under more realistic behavioral assumptions. A particularly promising direction comes from the demonstration in (Hart and Mas-Colell, 2000) that simple regret minimization learning leads to correlated equilibria rather than Nash equilibria. [^1] The Pluribus robot designed by researchers at Carnegie Mellon university, which beat the best professional human players at no-limit Texas hold ’em poker also uses a variation of regret minimization, suggesting that the algorithm is not just simple, but also capable. In addition, other adaptive procedures such as calibrated learning (due to (Foster and Vohra, 1997)) and universal conditional consistency (due to (Fudenberg and Levine, 1999)) have also been shown to approach correlated equilibria.222(Stoltz and Lugosi, 2007) demonstrate that regret minimization yields correlated equilibria when the strategy sets are convex and compact. These results suggests that correlated equilibrium may be a more natural solution concept for implementation problems, as it emerges from straightforward adaptive behavior rather than requiring sophisticated strategic reasoning.333(Halpern and Leung, 2016) and (Srivastava, 2019) provide empirical evidence in support of the claim that decision makers perform regret minimization.
While we know which social choice functions can be implemented in Nash equilibrium, we lack a characterization of SCFs which can be implemented in correlated equilibria and implementing mechanisms that explicitly target implementation in correlated equilibria. Such mechanisms would be especially valuable as they could achieve desired social outcomes under simple learning dynamics rather than requiring agents to compute equilibria.
This paper provides a step forward in resolving this tension between implementation theory and behavioral realism. We show that Maskin-monotonicity characterizes SCFs which are implementable in correlated equilibrium and provide a mechanism which can fully implement such SCFs as the outcome obtained in a unique correlated equilibrium. While there is a voluminous literature on full implementation, our paper offers the first mechanism that implements in adaptive heuristics per (Hart and Mas-Colell, 2000) . Previous papers such as (Cabrales, 1999) appeal to both integer games and adaptive dynamics where every best response must be played with positive probability, while (Cabrales and Serrano, 2011) requires that the agents switch with positive probability to any better or best response. In addition, we demonstrate in a simulation that the mechanism achieves the desirable social choice outcome in a canonical bilateral trade environment when the buyer and seller play the induced game using regret minimization heuristics. Finally, the mechanism does not rely on integer or modulo games, making it a reasonable candidate for practical deployment. [^4] See (Jackson, 1992) and (Moore, 1992) for a critique of these tail-chasing devices.
2. Mechanism and Proof of Implementation
2.1. Environment
Consider a finite set of agents $\mathcal{I}=\{1,2,...,I\}$ with $I\geq 2$; a finite set of possible states $\Theta$; and a set of pure alternatives $A$. We consider an environment with lotteries and transfers. Specifically, we work with the space of allocations/outcomes $X\equiv\Delta\left(A\right)\times\mathbb{R}^{I}$ where $\Delta(A)$ denotes the set of lotteries on $A$ that have a countable support, and $\mathbb{R}^{I}$ denotes the set of transfers to the agents.
Each state $\theta\in\Theta$ induces a type $\theta_{i}\in\Theta_{i}$ for each agent $i\in\mathcal{I}$. Assume that $\Theta$ has no redundancy, i.e., whenever $\theta\neq\theta^{\prime}$, we must have $\theta_{i}\neq\theta_{i}^{\prime}$ for some agent $i$. Hence, we can identify a state $\theta$ with its induced type profile $\left(\theta_{i}\right)_{i\in\mathcal{I}}$ and $\Theta$ with a subset of $\times_{i=1}^{I}\Theta_{i}$. Moreover, we say that a type profile $\left(\theta_{i}\right)_{i\in\mathcal{I}}$ identifies a state $\theta^{\prime}$ if $\theta_{i}=\theta_{i}^{\prime}$ for every $i\in\mathcal{I}$. Each type $\theta_{i}\in\Theta_{i}$ induces a utility function $u_{i}\left(\cdot,\theta_{i}\right):X\rightarrow\mathbb{R}$ which is quasilinear in transfers and has a bounded expected utility representation on $\Delta\left(A\right)$. That is, for each $x=\left(\ell,\left(t_{i}\right)_{i\in\mathcal{I}}\right)\in X$, we have $u_{i}\left(x,\theta_{i}\right)=v_{i}(\ell,\theta_{i})+t_{i}$ for some bounded expected utility function $v_{i}(\cdot,\theta_{i})$ over $\Delta\left(A\right)$. That is, we work with an environment with transferable utility (TU) restriction on agents’ preferences which is absent in (Maskin, 1999). As in (Abreu and Matsushima, 1992), we will take for granted that distinct elements of $\Theta_{i}$ induce different preference orderings over $\Delta\left(A\right),$ and also that a player is never indifferent over all elements of $A.$
We focus on a complete information environment in which the state $\theta$ is common knowledge among the agents but unknown to a mechanism designer. Thanks to the complete-information assumption, it is indeed without loss of generality to assume that agents’ values are private.
The designer’s objective is specified by a social choice function $f:\Theta\rightarrow X$, namely, if the state is $\theta$, the designer would like to implement the social outcome $f\left(\theta\right)$. We allow an SCF to be defined as a mapping from $\Theta$ to $X$ only so as to keep its consistency with the range of the outcome function used in the implementing mechanism. We can define $f:\Theta\rightarrow\Delta\left(A\right)$ as a special case of SCFs, as long as the designer is still allowed to impose off-the-equilibrium transfers in the implementing mechanism.
2.2. Implementation
We denote a mechanism by $\mathcal{M}=\left((M_{i},\tau_{i}\right)_{i\in\mathcal{I}},g)$ where $M_{i}$ is a nonempty set of messages available to agent $i$; $g:M\rightarrow X$ (where $M\equiv\times_{i=1}^{I}M_{i}$) is the outcome function; and $\tau_{i}:M\rightarrow\mathbb{R}$ is the transfer rule which specifies the payment to agent $i$. At each state $\theta\in\Theta$, the environment and the mechanism together constitute a game with complete information which we denote by $\Gamma(\mathcal{M},\theta)$.
Let $\sigma\in\Delta(M)$ be a probability distribution over $M$. A strategy profile $\sigma$ is said to be a correlated equilibrium of the game $\Gamma(\mathcal{M},\theta)$ if, for all agents $i\in\mathcal{I}$ and all messages $m_{i}$ such that $(m_{i},m_{-i})\insupp$$\left(\sigma\right)$ and $m_{i}^{\prime}\in M_{i}$, we have
$\sum_{m_{-i}\in M_{-i}}\sigma_{-i}(m_{-i}|m_{i})\left[\tilde{u}_{i}(g(m_{i},m_{-i}),\theta)+\tau_{i}(m_{i},m_{-i})\right]$
$\geq$ $\sum_{m_{-i}\in M_{-i}}\sigma_{-i}(m_{-i}|m_{i})\left[\tilde{u}_{i}(g(m_{i}^{\prime},m_{-i}),\theta)+\tau_{i}(m_{i}^{\prime},m_{-i})\right]\text{,}$
where we use $\sigma_{-i}(m_{-i}|m_{i})$ to denote the marginal probability on $m_{-i}$ conditional on $m_{i}$. For simplicity, we abuse notations to write $\sigma_{j}(m_{j}|m_{i})$ for the marginal probability on $m_{j}$ conditional on $m_{i}$.
Note that it is equivalent to have the following definition,
$\sum_{m_{-i}\in M_{-i}}\sigma(m_{i},m_{-i})\left[\tilde{u}_{i}(g(m_{i},m_{-i}),\theta)+\tau_{i}(m_{i},m_{-i})\right]$
$\geq$ $\sum_{m_{-i}\in M_{-i}}\sigma(m_{i},m_{-i})\left[\tilde{u}_{i}(g(m_{i}^{\prime},m_{-i}),\theta)+\tau_{i}(m_{i}^{\prime},m_{-i})\right]\text{.}$
Let $CE(\Gamma(\mathcal{M},\theta))$ denote the set of correlated equilibria of the game $\Gamma(\mathcal{M},\theta)$. We also denote by $\mbox{supp}\ (CE(\Gamma(\mathcal{M},\theta)))\$the set of message profiles that can be played with positive probability under some correlated equilibrium $\sigma\in CE(\Gamma(\mathcal{M},\theta)$, i.e.,
$\mbox{supp}\ (CE(\Gamma(\mathcal{M},\theta)))=\{m\in M:\mbox{there exists $\sigma\in CE(\Gamma(\mathcal{M},\theta))$ such that $\sigma(m)>0$}\}.$
We now define our notion of implementation.
Definition 2.1.
An SCF $f$ is implementable in correlated equilibria if there exists a mechanism $\mathcal{M}=\left((M_{i},\tau_{i}\right)_{i\in\mathcal{I}},g)$ such that for every state $\theta\in\Theta$, $m\in\mbox{supp}\ (CE(\Gamma(\mathcal{M},\theta)))\Rightarrow g\left(m\right)=f\left(\theta\right)$ and $\tau_{i}\left(m\right)=0$ for every $i\in\mathcal{I}$.
2.3. Maskin Monotonicity
For $\left(x,\theta_{i}\right)\in X\times\Theta_{i}$, we use $\mathcal{L}_{i}\left(x,\theta_{i}\right)$ to denote the lower-contour set at allocation $x$ in $X$ for type $\theta_{i}$, i.e.,
$\mathcal{L}_{i}\left(x,\theta_{i}\right)=\left\{x^{\prime}\in X:u_{i}\left(x,\theta_{i}\right)\geq u_{i}(x^{\prime},\theta_{i})\right\}.$
We use $\mathcal{SU}_{i}\left(x,\theta_{i}\right)$ to denote the strict upper-contour set of $x\in X$ for type $\theta_{i}$, i.e.,
$\mathcal{SU}_{i}\left(x,\theta_{i}\right)=\left\{x^{\prime}\in X:u_{i}(x^{\prime},\theta_{i})>u_{i}(x,\theta_{i})\right\}\text{.}$
We now state the definition of Maskin monotonicity which (Maskin, 1999) proposes for Nash implementation.
Definition 2.2.
An SCF $f$ satisfies Maskin monotonicity if, for every pair of states $\tilde{\theta}$ and $\theta$ with $f(\tilde{\theta})\not=f\left(\theta\right)$, there is some agent $i\in\mathcal{I}$ such that
(1) $\mathcal{L}_{i}(f(\tilde{\theta}),\tilde{\theta}_{i})\cap\mathcal{SU}_{i}(f(\tilde{\theta}),\theta_{i})\not$$\varnothing\text{.}$
The agent $i$ in Definition 2.2 is called a “whistle-blower” or a “test agent”; likewise, an allocation in $\mathcal{L}_{i}(f(\tilde{\theta}),\tilde{\theta}_{i})\cap\mathcal{SU}_{i}(f(\tilde{\theta}),\theta_{i})$ is called a “test allocation” for agent $i$ and the ordered pair of states $(\tilde{\theta},\theta)$.
2.4. The Mechanism
In this section, we present our main result which shows that Maskin monotonicity is necessary and sufficient for implementation in correlated equilibrium. We formally state the result as follows:
Theorem 2.3.
An SCF $f$ is implementable in correlated equilibria if and only if it satisfies Maskin monotonicity.
(Maskin, 1999) establishes that Maskin monotonicity is necessary for implementation in Nash equilibria. Since each Nash equilibrium is also a correlated equilibrium, Maskin monotonicity is also necessary for implementation in correlated equilibria. To show that it is sufficient, we will construct an implementing mechanism in the remainder of the section having first established some preliminaries.
2.4.1. Dictator Lotteries
Lemma 0.
For each agent $i\in\mathcal{I}$, there exists a function $y_{i}:\Theta_{i}\rightarrow X$ such that for every types $\theta_{i}$ and $\theta_{i}^{\prime}$ with $\theta_{i}\neq\theta_{i}^{\prime}$, we have
(2) $u_{i}\left(y_{i}\left(\theta_{i}\right),\theta_{i}\right)>u_{i}\left(y_{i}\left(\theta_{i}^{\prime}\right),\theta_{i}\right)\text{.}$
(Abreu and Matsushima, 1992) prove the existence of lotteries $\left\{y_{i}\left(\cdot\right)\right\}\subset\Delta\left(A\right)$ which satisfy Condition (2). We now define a notion called the best challenge scheme, which plays a crucial role in proving Theorem 2.3. First, a challenge scheme for an SCF $f$ is a collection of (pre-assigned) test allocations $\{x(\tilde{\theta},\theta_{i})\}$, one for each pair of state $\tilde{\theta}$ and type $\theta_{i}$ of agent $i$, such that
(3) $\text{ if }\mathcal{L}_{i}(f(\tilde{\theta}),\tilde{\theta}_{i})\cap\mathcal{SU}_{i}(f(\tilde{\theta}),\theta_{i})$ $\neq$ , $\varnothing\text{, then }x(\tilde{\theta},\theta_{i})\in\mathcal{L}_{i}(f(\tilde{\theta}),\tilde{\theta}_{i})\cap\mathcal{SU}_{i}(f(\tilde{\theta}),\theta_{i})\text{;}$
[…]
When $\mathcal{L}_{i}(f(\tilde{\theta}),\tilde{\theta}_{i})\cap\mathcal{SU}_{i}(f(\tilde{\theta}),\theta_{i})\neq\varnothing$, we may think of state $\tilde{\theta}$ as an announcement made by another agent(s) which agent $i$ could challenge (as a whistle-blower) when agent $i$ has true type $\theta_{i}$. The following lemma shows that there is a challenge scheme in which each whistle-blower $i$ facing state announcement $\tilde{\theta}$ has a weak incentive to report his true type $\theta_{i}$ to challenge $\tilde{\theta}$.
Lemma 0.
For any SCF $f$, there is a challenge scheme $\{x(\tilde{\theta},\theta_{i})\}_{i\in\mathcal{I},\tilde{\theta}\in\Theta,\theta_{i}\in\Theta_{i}}$ such that for every $i\in\mathcal{I},\ \tilde{\theta}\in\Theta,$ and $\theta_{i}\in\Theta_{i}$,
(5) $u_{i}(x(\tilde{\theta},\theta_{i}),\theta_{i})\geq u_{i}(x(\tilde{\theta},\theta_{i}^{\prime}),\theta_{i}),\forall\theta_{i}^{\prime}\in\Theta_{i}\text{;}$
Having formulated the best challenge scheme, we will detail the mechanism in the following sections.
2.4.2. Message Space and Outcome Function
For each agent $i$, we define the message space as follows. A generic message
$m_{i}=(m_{i}^{1},m_{i}^{2})\in M_{i}^{1}\times M_{i}^{2},$
where $M_{i}^{1}=M_{i}^{2}=\times_{j=1}^{I}\Theta_{j}$ is the state space out of which each agent reports messages twice.
For each message profile $m\in M$, the allocation is determined as follows:
$g\left(m\right)=\frac{1}{I^{2}}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}g_{i,j}(m)\text{,}$
and,
$g_{i,j}(m)=\left[e_{i,j}\left(m_{i},m_{j}\right)\left(\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}(m_{k,k}^{1})\right)\oplus\left(1-e_{i,j}\left(m_{i},m_{j}\right)\right)x(m_{i}^{2},m_{j,j}^{1})\right]$
where $\{y_{k}(\cdot)\}\$are the dictator lotteries for agent $k$ obtained from Lemma 2.4, moreover, we define
$e_{i,j}(m_{i},m_{j})=\left\{\begin{array}[]{ll}0\text{,}&\text{if }m_{i}^{2}\in\Theta\text{, and }x(m_{i}^{2},m_{j,j}^{1})=f(m_{i}^{2})\text{;}\\ \varepsilon\text{,}&\text{if }m_{i}^{2}\in\Theta\text{, and }x(m_{i}^{2},m_{j,j}^{1})\not=f(m_{i}^{2})\text{;}\\ 1\text{,}&\text{if }m_{i}^{2}\notin\Theta\text{.}\end{array}\right$.
Hereafter we say agent $i$ is challenged by agent $j$ if $x(m_{i}^{2},m_{j,j}^{1})\not=f(m_{i}^{2})$, which is equivalent to $g_{i,j}\not=f(m_{i}^{2})$.
In words, the designer first chooses an ordered pair(including agent $i$ paired with himself) of agents $(i,j)$ with equal probability. The outcome function distinguishes three cases: (1) if agent $i$ reports a type profile which identifies a state in $\Theta,$ and agent $j$ does not challenge agent $i$, then we implement
$g_{i,j}(m)=f\left(m_{i}^{2}\right);$
(2) if agent $i$ reports a type profile which identifies a state in $\Theta,$ and agent $j$ does challenge agent $i$, then we implement the compound lottery [^5] More precisely, if $x=\left(\ell,\left(t_{i}\right)_{i\in\mathcal{I}}\right)$ and $x^{\prime}=\left(\ell^{\prime},\left(t_{i}^{\prime}\right)_{i\in\mathcal{I}}\right)$ are two outcomes in $X$, we identify $\alpha x\oplus\left(1-\alpha\right)x^{\prime}$ with the outcome $\left(\alpha\ell\oplus\left(1-\alpha\right)\ell^{\prime},\left(\alpha t_{i}+\left(1-\alpha\right)t_{i}^{\prime}\right)_{i\in\mathcal{I}}\right)$. For simplicity, we also write the compound lottery as $\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}(m_{k}^{1})$.
$g_{i,j}(m)=C_{i,j}^{\varepsilon}(m_{i},m_{j})\equiv\varepsilon\left(\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}\left(m_{k,k}^{1}\right)\right)\oplus\left(1-\varepsilon\right)x(m_{i}^{2},m_{j,j}^{1})\text{.}$
Note that $C_{i,j}^{\varepsilon}(m_{i},m_{j})$ is an $\left(\varepsilon,1-\varepsilon\right)$-combination of (i) the dictator lotteries which occur with equal probability; and (ii) the allocation specified by the best challenge scheme $x(m_{i}^{2},m_{j,j}^{1})$;
(3) if agent $i$ reports a type profile which does not identify a state in $\Theta$, then we implement the dictator lottery $\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}\left(m_{k,k}^{1}\right).$
We abuse notations to write $x(m_{i}^{2},m_{j,j}^{1})\$and $f(m_{i}^{2})$ for all $i$ and $j$ assuming $m_{i}^{2}$ identifies a state since the outcome will be only determined by the dictator lottery otherwise.
Given the construction of dictator lotteries, we can choose $\varepsilon>0$ sufficiently small, and $\eta>0$ sufficiently large such that firstly, we have
(6) $\eta>\sup_{i\in\mathcal{I},\theta_{i}\in\Theta_{i},m,m^{\prime}\in M}\left|u_{i}(g\left(m\right),\theta_{i})-u_{i}(g\left(m^{\prime}\right),\theta_{i})\right|\text{;}$
secondly, it does not disturb the “effectiveness” of agent $j$’s challenge: due to (3), we can have
$x(m_{i}^{2},m_{j,j}^{1})$ $\neq$ $f(m_{i}^{2})\Rightarrow$
(7) $u_{j}(C_{i,j}^{\varepsilon}(m_{i},m_{j}),m_{i,j}^{2})$ $<$ $u_{j}(f(m_{i}^{2}),m_{i,j}^{2})\text{ and }u_{j}(C_{i,j}^{\varepsilon}(m_{i},m_{j}),m_{j,j}^{1})>u_{j}(f(m_{i}^{2}),m_{j,j}^{1})\text{.}$
It means that whenever agent $j$ challenges agent $i$, the lottery $C_{i,j}^{\varepsilon}(m_{i},m_{j})$ is strictly worse than $f\left(m_{i}^{2}\right)$ for agent $j$ when agent $i$ tells the truth about agent $j$’s preference in $m_{i}^{2}$; moreover, the lottery $C_{i,j}^{\varepsilon}(m_{i},m_{j})$ is strictly better than $f\left(m_{i}^{2}\right)$ for agent $j$ when agent $j$ tells the truth in $m_{j,j}^{1}$, which implies that agent $i$ tells a lie about agent $j$’s preference.
2.4.3. Transfer Rule
We now define the transfer rule. For every message profile $m\in M$ and every agent $i\in\mathcal{I}$, we specify the transfer received by agent $i$ as follows:
$\tau_{i}(m)=\sum_{j\neq i}\left[\tau_{i,j}^{1}(m_{i},m_{j})+\tau_{i,j}^{2}(m_{i},m_{j})+\tau_{i,j}^{3}(m_{i},m_{j})\right].$
(10) $\tau_{i,j}^{1}\left(m_{i},m_{j}\right)$ $=$ $\left\{\begin{tabular}[]{ll}$-2\eta$ ,&\text{if} $x(m_{i}^{2},m_{j,j}^{1})\neq f(m_{i}^{2})$;\\ $0$,&\text{otherwise.}\end{tabular}\ \right$.
[…]
(17) $\tau_{i,j}^{3}\left(m_{i},m_{j}\right)$ $=$ $\left\{\begin{array}[]{ll}-\eta\text{,}&\text{if }x(m_{i}^{2},m_{j,i}^{1})\neq f(m_{i}^{2}){;}\\ 0\text{,}&\text{otherwise.}\end{array}\right$.
Recall that $\eta>0$ is chosen to be greater than the maximal utility difference from the outcome function $g\left(\cdot\right)$; see (6), and $\epsilon$ is an arbitrarily small positive number.
Given each agent $j\neq i$, we define the first transfer $\tau^{1}$ for agent $i$ so that agent $i$ is asked to pay $2\eta$ if agent $i$ is challenged by agent $j$ or agent $i$ reports a type profile which does not identify a state in $\Theta$.
We will first show that in equilibrium, no agent $i$ reports $m_{i}^{2}$ which does not identify a state in $\Theta$ because it immediately incurs the largest penalty for agent $i$ regarding how to report $m_{i}^{2}.$
For $\tau^{2},$ it is designed so that agent $i$ is asked to pay $\epsilon$ if agent $j$ challenges himself, and agent $i$’s first report on agent $j$’s type differs from the reported one from agent $j$ in agent $j$’s first report. Note that later we will argue that whenever agent $j$ challenges himself agent $j$ reports the truth in his own type from the first report, thus, it is a weakly dominant strategy for agent $i$ to report agent $j$’s true type in agent $i$’s first report.
For $\tau^{3},$ it is designed such that given agent $j$ makes a positive bet on agent $i$ challenging himself, agent $i$ is asked to pay $\eta$ if his second report on his own type is different from the one reported by agent $j$ in agent $j$’s first report.
2.5. Proof of Implementation
We outline our proof strategy as follows. We will first show that in equilibrium, no agent $i$ reports $m_{i}^{2}$ which does not identify a state in $\Theta.$ Then, we prove that every agent reports a state which induces the socially desired outcome and no one challenges anyone, thus, there is no transfer incurred on equilibrium path.
  1. (1)
    First, we show that in equilibrium, if an agent $i$ reports a state in his second report, inducing an outcome different from the socially desired one (we call it a false state), then agent $i$ must be challenged (maybe by himself). This is established by monotonicity.
  2. (2)
    Second, in equilibrium, no one reports a false state challenged by another agent. This is obtained by the first step and the penalty imposed on agent if he is challenged.
  3. (3)
    Third, we show that no one reports a false state challenged by himself. This is the key step and difficult point in the proof.
    1. (a)
      We first show that if agent $j$ has a positive belief that agent $i$ challenges himself, then agent $j$ must report agent $i$’s true type in agent $j$’s first report (Note that this is due to $\tau^{2}$ and the scale of $\epsilon$ can be arbitrarily small since agent $j$’s first report about agent $i$’s type only affect agent $i$ through $\tau_{i}^{2}$).
    2. (b)
      Second, we show that from the agent $i$’s perspective, if he challenges himself, then all the messages from his opponent report the agent $i$’s true type in their first reports, thus, agent $i$ suffers $\eta$ with probability $1$ according to $\tau_{i}^{3}$.
    3. (c)
      The third transfer then provides a strict incentive for agent $i$ to report the truth in his second report instead of reporting a lie being challenged by himself.
    We then show that in the equilibrium path, the agent $i$ does not challenge himself.
Now, for each state $\tilde{\theta}\in\Theta$, we define the following set of agents:
$\mathcal{J}(\tilde{\theta})\equiv\left\{j\in\mathcal{I}:\mathcal{L}_{j}(f(\tilde{\theta}),\tilde{\theta}_{j})\cap\mathcal{SU}_{j}(f(\tilde{\theta}),\theta_{j})=\varnothing\right\}\text{.}$
Claim 1.
Given an arbitrary agent $i$ and message profile $m\insupp$$\left(\sigma\right)$, we have $x(m_{i}^{2},m_{j,j}^{1})\not=f(m_{i}^{2})\$if and only if $j\not\in\mathcal{J}\left(m_{i}^{2}\right)$.
Proof.
Fix agent $i\in\mathcal{I}$ and a message profile $m\insupp$$(\sigma)$. We first prove the if-part. Suppose, on the contrary, that $x(m_{i}^{2},m_{j,j}^{1})=f(m_{i}^{2})$. Then, consider the deviation from $m_{j}$ to $\tilde{m}_{j}$ such that $\tilde{m}_{j,j}^{1}=\theta_{j}$ while $m_{j}$ and $\tilde{m}_{j}$ are the same in every other dimension. By (7), it delivers a strictly better payoff for agent $j$ against $m_{-j}$ where $j\not\in\mathcal{J}\left(m_{i}^{2}\right)$. At the same time, by Lemmas 2.4 and 2.5, the deviation from $m_{i}$ to $\tilde{m}_{i}$ generates no payoff loss for agent $i$ against any $m_{-i}^{\prime}\not=m_{-i}$. Thus, the deviation $\tilde{m}_{j}$ is profitable, which contradicts the hypothesis that $\sigma$ is an equilibrium of the game $\Gamma(\mathcal{M},\theta)$.
Next, we prove the only-if-part. Suppose, on the contrary, that there exists some agent $j\in\mathcal{J}\left(m_{i}^{2}\right)$ such that $x(m_{i}^{2},m_{j,j}^{1})\not=f(m_{i}^{2}).$ Since $j\in\mathcal{J}\left(m_{i}^{2}\right)$, we must have $m_{j,j}^{1}\neq\theta_{j}$. Define $\tilde{m}_{j}$ as a deviation which is identical to $m_{j}$ except that $\tilde{m}_{j,j}^{1}=\theta_{j}\neq m_{j,j}^{1}$. Then we have $x(m_{i}^{2},\tilde{m}_{j,j}^{1})=x(m_{i}^{2},\theta_{j})=m_{i}^{2}$ since $j\in\mathcal{J}\left(m_{i}^{2}\right).$ By (7), $\tilde{m}_{j}$ generates a strictly better payoff for agent $j$ than $m_{j}$ against $m_{-j}$. By Lemmas 2.4 and 2.5, we also know that agent $j$’s payoff generated by $\tilde{m}_{j}$ is at least as good as that generated by $m_{j}$ against any $m_{-j}^{\prime}\neq m_{-j}$. Hence, $\tilde{m}_{j}$ constitutes a profitable deviation, which contradicts the hypothesis that $\sigma$ is an equilibrium of the game $\Gamma(\mathcal{M},\theta)$. ∎
Claim 2.
No one challenges an allocation announced in the second report of any other agent, i.e., for any pair of agents $i,j\in\mathcal{I}$ with $i\neq j$ and any $m\insupp$$\left(\sigma\right)$, $x(m_{i}^{2},m_{j,j}^{1})=f(m_{i}^{2})$.
Proof.
Suppose to the contrary that there exist $i,j\in\mathcal{I}$ with $i\neq j$, $m\in\mbox{supp}(\sigma)$ such that $x(m_{i}^{2},m_{j,j}^{1})\neq f(m_{i}^{2})$. By Claim 1, $j\not\in\mathcal{J}\left(m_{i}^{2}\right)$. By Claim 1, we know that for any agent $j\not\in\mathcal{J}\left(m_{i}^{2}\right),$ we have $x(f(m_{i}^{2}),\hat{m}_{j,j}^{1})\neq f(m_{i}^{2})$, for every $\hat{m}_{-i}$ such that $\sigma(m_{i},\hat{m}_{-i})>0$. In addition, $\hat{m}_{j}^{1}=\theta_{j}$ for every $j$ such that $j\not\in\mathcal{J}\left(m_{i}^{2}\right).$ Hence, from agent $i$’s perspective, conditional on playing $m_{i},$ he is challenged with probability $1.$ Thus, agent $i$ is penalized $2\eta$ due to (10)). Consider a deviation to $\tilde{m}_{i}$ which is the same as $m_{i}$ except that $\tilde{m}_{i}^{2}=\left(\theta_{j}\right)_{j\in\mathcal{I}}$. Note that from agent $i$’s perspective, dictator lotteries are triggered with probability one, hence, all the agents report the truth in the first reports. Then, agent $i$ avoids paying the penalty $2\eta$ for being challenged, while the potential loss from allocation is bounded by $\eta,$ the loss from (17) is bounded by $\eta.$ Therefore, it is a profitable deviation. This contradicts to that $\sigma$ is an equilibrium. ∎
Claim 3.
For any pair of agents $i$ and $j$ such that $i\neq j$, message profile $m\insupp$$\left(\sigma\right)$, if $x(m_{j}^{2},m_{j,j}^{1})\neq f(m_{j}^{2})$, then we have $m_{i,j}^{1}=\theta_{j}$ and $m_{i,i}^{1}=\theta_{i}$.
Proof.
First, we fix an arbitrary correlated equilibrium $\sigma$. Given $m_{i}$ we collect all the opponents’ message profile together with $m_{i}$ agent $i$ knows that agent $j$ is challenged by himself:
$E_{j}(m_{i})=\{\tilde{m}_{-i}:x(\tilde{m}_{j}^{2},\tilde{m}_{j,j}^{1})\neq f(\tilde{m}_{j}^{2})\text{ and }\sigma(m_{i},\tilde{m}_{-i})>0.\}$
From the condition in Claim 3, we know that $E_{j}(m_{i})\neq\varnothing$. Conditional on $m_{i}$, dictator lotteries are triggered with positive probability; hence by Lemmas 2.4 and 2.5, $m_{i,i}^{1}=\theta_{i}$. Note that for any $\tilde{m}_{-i}$ such that $\sigma_{i}(\tilde{m}_{-i}|m_{i})>0$ and $\tilde{m}_{-i}\not\in E_{j}(m_{i}),$ $j$ does not challenge himself. Thus by Claim 1, $x(\tilde{m}_{j}^{2},\theta_{j})=f(\tilde{m}_{j}^{2}).$ Hence, according to $\tau_{i,j}^{2}$ (see (14)), we have $m_{i,j}^{1}=\theta_{j}.$ In addition, when agent $j$ is challenged, we have $m_{j,j}^{1}=\theta_{j}$.∎
Claim 4.
For every agent $j$, message profile $m\insupp$$\left(\sigma\right)$, we have $x(m^{2}_{j},m^{1}_{k,k})=f(m^{2}_{j})$ for every agent $k$.
Proof.
By Claim 2, we know that for every $k\neq j$, Claim 4 holds. It remains to show that Claim 4 holds for $k=j$. Suppose there exists $m\insupp$$\left(\sigma\right)$, such that $x(m_{j}^{2},m_{j,j}^{1})\not=f(m_{j}^{2})$. By Lemmas 2.4 and 2.5, $m_{j,j}^{1}=\theta_{j}$. Now we show that for agent $j$ who reports $m_{j}$ with $m_{j,j}^{2}\not=\theta_{j}$, it is strictly better for agent $j$ to deviate to $\tilde{m}_{j}$ which is identical to $m_{j}$ but $\tilde{m}_{j,j}^{2}=\theta$ the true type profile. Specifically, due to Claim 3, for every $\tilde{m}_{-j}$ such that $\sigma(m_{j},\tilde{m}_{-j})>0$, for every agent $i\neq j$, we have $\tilde{m}_{i,j}^{3}>0$, $\tilde{m}_{i,j}^{1}=\theta_{j}$ and $\tilde{m}_{i,i}^{1}=\theta_{i}$. By $\tau^{3}$ in (17), it is strictly better, and for $\tau^{1}$ and $\tau^{2},$ there is no loss incured. Hence, we know that the deviation is profitable. Thus, it contradicts the hypothesis that $x(m_{j}^{2},m_{j,j}^{1})\not=f(m_{j}^{2})$. ∎
Claim 5.
For every $m\in\text{supp}(\sigma)$, $g(m)=f(\theta)$, and $\tau_{i,j}(m)=0$ for every agent $i$ and $j$.
Proof.
By Claim 1 and Claim 4, we know that for every $m\in\text{supp}(\sigma)$, $g(m)=f(\theta)$, and $\tau_{i,j}^{1}=0$ for every agent $i$ and $j$. Due to the construction of the proper scoring rule $sc$, we know that $m_{i,j}^{3}=0$ for every agent $i$ and $j$. Hence $\tau_{i,j}^{2}=\tau_{i,j}^{3}=0$ for every agent $i$ and $j$. Hence, we achieve implementation with no transfers incurred. ∎
Remark 1.
Throughout the proof of our main theorem, to argue the behavior on the equilibrium path, whenever we use a possible profitable deviation to derive a contradiction, we use the joint true type profile. Therefore, a feature of our mechanism is that even with a restricted message space, as long as the true type profiles are available, our mechanism would still implement.
2.6. Social Choice Correspondences
A large portion of the implementation literature strives to deal with social choice correspondences (hereafter, SCCs), i.e., multi-valued social choice rules. In this section, we extend our implementation result to cover the case of SCCs. We suppose that the designer’s objective is specified by an SCC $F:\Theta\rightrightarrows X$; and for simplicity, we assume that $F\left(\theta\right)$ is a finite set for each state $\theta\in\Theta$. It includes the special case where the co-domain of $F$ is $A$. Following (Maskin, 1999), we first define the notion of Nash implementation for an SCC.
Definition 2.6.
An SCC $F$ is implementable in correlated equilibria by a finite mechanism if there exists a mechanism $\mathcal{M}=\left((M_{i},\tau_{i})_{i\in\mathcal{I}},g\right)$ such that for every state $\theta\in\Theta$, the following two conditions are satisfied: (i) for every $x\in F(\theta)$, there exists a pure-strategy Nash equilibrium mm\ in the game $\Gamma(\mathcal{M},\theta)$ with $g(m)=x$ and $\tau_{i}(m)=0$ for every agent $i\in\mathcal{I}$; and (ii) for every $m\in\mbox{supp}\ (CE(\Gamma(\mathcal{M},\theta)))$, we have $\mbox{supp}(g(m))\subseteq F\left(\theta\right)$ and $\tau_{i}\left(m\right)=0$ for every agent $i\in\mathcal{I}$.
Second, we state the definition of Maskin monotonicity for an SCC.
Definition 2.7.
[…]
We now state our implementation result for SCCs and relegate the proof to Appendix 2.8.666When there are only two agents, we can still show that every Maskin-monotonic SCC $F$ is weakly implementable in Nash equilibria, that is, there exists a mechanism which has a pure-strategy Nash equilibrium and satisfies requirement (ii) in Definition 2.6.
Theorem 2.8.
Suppose there are at least three agents. An SCC $F$ is implementable in correlated equilibria by a finite mechanism if and only if it satisfies Maskin monotonicity.
[…]
Remark. Discuss NK here.
2.7. Proof of Theorem 2.8
[…]
As in the case of SCFs, the following lemma shows that there is a challenge scheme under which truth-telling induces the best allocation.
Lemma 0.
[…]
(18) $u_{i}(x(\tilde{\theta},z,\theta_{i}),\theta_{i})\geq u_{i}(x(\tilde{\theta},z,\theta_{i}^{\prime}),\theta_{i}),\forall\theta_{i}^{\prime}\in\Theta_{i}\text{;}$
Lemma 18 is established when we apply the proof of Lemma 2.5 to the challenge scheme $\{x(\tilde{\theta},z,\theta_{i})\}_{i\in\mathcal{I},\tilde{\theta}\in\Theta,z\in F(\tilde{\theta}),\theta_{i}\in\Theta_{i}}$. Thus, we omit the proof here.
Next, we propose a mechanism $\mathcal{M}=\left((M_{i}),g,(\tau_{i})\right)_{i\in\mathcal{I}}$ which will be used to prove the if-part of Theorem 2.8. First, a generic message of agent $i$ is described as follows:
$m_{i}=\left(m_{i}^{1},m_{i}^{2},m_{i}^{3},m_{i}^{4}\right)\in M_{i}=M_{i}^{1}\times M_{i}^{2}\times M_{i}^{3}\times M_{i}^{4}=\Theta\times\Theta\times F(\Theta)\times C_{i}^{1}\text{ s.t.}$
$m_{i}^{3}\in F(m_{i}^{2})\text{. }$
That is, agent $i$ is asked to announce (1) agent $i$’s own type (which we denote by $m_{i}^{1}$); (2) a type profile (which we denote by $m_{i}^{2}$); (3) an allocation $m_{i}^{3}$ such that $m_{i}^{3}\in F(m_{i}^{2})$ if $m_{i}^{2}$ is a state;(4) $C_{i}^{1}=\times_{j\neq i}C_{i,j}$ with $C_{i,j}=[0,1]$, where each agent reports the information about his opponents’ status on “being challenged”(defined as follows). As we do in the case of SCFs, we write $m_{i,j}^{2}=\tilde{\theta}_{j}$ if agent $i$ reports in $m_{i}^{2}$ that agent $j$’s type is $\tilde{\theta}_{j}$.
We define $\phi(m)$ as follows: for each $m\in M$,
ϕ(m)={
x,
m13,
if |{iI:=mi3x}|-I1;
otherwise.
[…]
We say that $\phi(m)$ is an effective allocation under $m$. In words, the effective allocation is $x$, if there are I1I-1 players who agree on allocation $x$; otherwise, the effective allocation is the allocation announced by agent 1.
The allocation rule $g$ is defined as follows: for each $m\in M$,
$g\left(m\right)=\frac{1}{I^{2}}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}g_{i,j}(m)\text{,}$
and,
$g_{i,j}(m)=\left[e_{i,j}\left(m_{i},m_{j}\right)\left(\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}(m_{k,k}^{1})\right)\oplus\left(1-e_{i,j}\left(m_{i},m_{j}\right)\right)x(\tilde{\theta},\phi(m),m_{j,j}^{1})\right].$
where $\{y_{k}(\theta_{k})\}_{\theta_{k}\in\Theta_{k}}$ are the dictator lotteries for agent $k$ as defined in Lemma 2.4. Given a message profile $m,$ and a pair of agents $i$ and $j,$ we say that agent $j$ challenges agent $i$ if and only if $m_{i}^{3}=\phi(m)$ and $x(m_{i}^{2},\phi(m),m_{j,j}^{1})\neq\phi(m)$, i.e., agent $i$’s reported allocation is an effective one and agent $j$ challenges this effective allocation. We define the $e_{i,j}$-function as follows: for each $m\in M$,
$e_{i,j}(m_{i},m_{j})=\left\{\begin{array}[]{ll}\varepsilon\text{,}&\text{if agent }j\text{ challenges agent }i\text{;}\\ 0\text{,}&\text{otherwise }.\end{array}\right$.
Recall that the $e_{i,j}$-function in Section 2.4.2 for the case of SCFs only depends on $m_{i}$ and $m_{j}$. In contrast, the $e_{i,j}$-function here depends on the entire message profile, as the nature of the challenge also depends on whether the allocation reported by agent $i$ is an effective allocation or not.
Fix $i,j\in\mathcal{I},\varepsilon\in(0,1)$, and $m\in M$. Then, we define
$C_{i,j}^{\varepsilon}(m)\equiv\varepsilon\times\frac{1}{I}\sum_{k\in\mathcal{I}}y_{k}\left(m_{k,k}^{1}\right)\oplus\left(1-\varepsilon\right)\times x(m_{i}^{2},\phi(m),m_{j,j}^{1})\text{.}$
For every message profile $m$ and agent $j$, we can choose $\varepsilon>0$ sufficiently small such that (i) $C_{i,j}^{\varepsilon}(m)$ does not disturb the “effectiveness” of agent $j$’s challenge, i.e.,
$x(m_{i}^{2},\phi(m),m_{j}^{1})$ $\neq$ $\phi(m)\Rightarrow$
(19) $u_{j}(C_{i,j}^{\varepsilon}(m),m_{i,j}^{2})$ $<$ $u_{j}(\phi(m),m_{i,j}^{2})\text{ and }u_{j}(C_{i,j}^{\varepsilon}(m),m_{j,j}^{1})>u_{j}(\phi(m),m_{j,j}^{1})\text{.}$
2.7.1. Transfer Rule
We now define the transfer rule. For every message profile $m\in M$ and every agent $i\in\mathcal{I}$, we specify the transfer received by agent $i$ as follows:
$\tau_{i}(m)=\sum_{j\neq i}\left[sc_{i,j}(m)+\tau_{i,j}^{1}(m_{i},m_{j})+\tau_{i,j}^{2}(m_{i},m_{j})+\tau_{i,j}^{3}(m_{i},m_{j})\right].$
For the transfer $sc_{i,j}$, we define it as a strict scoring rule, a “bet” on his opponent agent $j$ challenging agent $j$ himself. Thus, the transfers depend on the probability reported by agent $i$ on the event agent $j$ being challenged, $c_{i,j}$ and the one realized. We define the proper scoring rule explicitly as follows,
$sc_{i,j}\left(m\right)=\left\{\begin{array}[]{ll}-(c_{i,j})^{2}-(1-c_{i,j})^{2}+2c_{i,j}-1\text{, }&\text{if agent }j\text{ challenges himself};\\ -(c_{i,j})^{2}-(1-c_{i,j})^{2}+2(1-c_{i,j})-1\text{,}&\text{otherwise.}\end{array}\right$.
(22) $\tau_{i,j}^{1}\left(m_{i},m_{j}\right)$ $=$ $\left\{\begin{tabular}[]{ll}$-2\eta$ ,&\text{if} $\text{agent }j\text{ challenges agent }i$;\\ $0$,&\text{otherwise.}\end{tabular}\ \right$.
[…]
(28) $\tau_{i,j}^{3}\left(m_{i},m_{j}\right)$ $=$ $\left\{\begin{array}[]{ll}-\eta\text{,}&\text{if }c_{j,i}>0\text{ and }m_{i,i}^{2}\not=m_{j,i}^{1}{;}\\ 0\text{,}&\text{otherwise.}\end{array}\right$.
Recall that $\eta>0$ is chosen to be greater than the maximal utility difference from the outcome function $g\left(\cdot\right),$ and $\varepsilon$ is a small positive nubmer; see (6).
In the rest of the proof , we fix $\theta$ as the true state and $\sigma$ as a correlated equilibrium of the game $\Gamma(\mathcal{M},\theta)$ throughout.
2.7.2. Proof of Implementation
[…]
$\mathcal{J}\left(\tilde{\theta},x\right)\equiv\left\{j\in\mathcal{I}:\mathcal{L}_{j}(x,\tilde{\theta}_{j})\cap\mathcal{SU}_{j}(x,\theta_{j})=\varnothing\right\}\text{.}$
Claim 6.
Given an arbitrary agent $i$ and message profile $m\insupp$$\left(\sigma\right)$ such that $m_{i}^{3}=\phi(m)$, we have $x(m_{i}^{2},m_{i}^{3},m_{j,j}^{1})\not=m_{i}^{3}\$if and only if $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$.
Proof.
Fix agent $i\in\mathcal{I}$ and a message profile $m\insupp$$(\sigma)$. We first prove the if-part. Suppose, on the contrary, that there exists some agent $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$ such that $x(m_{i}^{2},m_{i}^{3},m_{j,j}^{1})=m_{i}^{3}$. Then, consider the deviation from $m_{j}$ to $\tilde{m}_{j}$ such that $\tilde{m}_{j,j}^{1}=\theta_{j}$ while $m_{j}$ and $\tilde{m}_{j}$ are the same in every other dimension. By (19), it delivers a strictly better payoff for agent $j$ against $m_{-j}$ where $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$. At the same time, by Lemmas 2.4 and 2.9, the deviation from $m_{j}$ to $\tilde{m}_{j}$ generates no payoff loss for agent $j$ against any $m_{-j}^{\prime}\not=m_{-j}$. Thus, the deviation $\tilde{m}_{j}$ is profitable, which contradicts the hypothesis that $\sigma$ is an equilibrium of the game $\Gamma(\mathcal{M},\theta)$.
Next, we prove the only-if-part. Suppose, on the contrary, that there exists some agent $j\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$ such that $x(m_{i}^{2},m_{i}^{3},m_{j,j}^{1})\not=m_{i}^{3}.$ Since $j\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$, we must have $m_{j,j}^{1}\neq\theta_{j}$. Define $\tilde{m}_{j}$ as a deviation which is identical to $m_{j}$ except that $\tilde{m}_{j,j}^{1}=\theta_{j}\neq m_{j,j}^{1}$. Then we have $x(m_{i}^{2},m_{i}^{3},\theta_{j})=m_{i}^{3}$ since $j\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right).$ By (19), $\tilde{m}_{j}$ generates a strictly better payoff for agent $j$ than $m_{j}$ against $m_{-j}$. By Lemmas 2.4 and 2.9, we also know that agent $j$’s payoff generated by $\tilde{m}_{j}$ is at least as good as that generated by $m_{j}$ against any $m_{-j}^{\prime}\neq m_{-j}$. Hence, $\tilde{m}_{j}$ constitutes a profitable deviation, which contradicts the hypothesis that $\sigma$ is an equilibrium of the game $\Gamma(\mathcal{M},\theta)$. ∎
Claim 7.
No one challenges an allocation announced by any other agent, i.e., for any pair of agents $i,j\in\mathcal{I}$ with $i\neq j$ and any $m\insupp$$\left(\sigma\right)$ such that $m_{i}^{3}=\phi(m)$, $x(m_{i}^{2},m_{i}^{3},m_{j,j}^{1})=m_{i}^{3}$.
Proof.
Suppose to the contrary that there exist $i,j\in\mathcal{I}$ with $i\neq j$, $m\in\mbox{supp}(\sigma)$ such that such that $m_{i}^{3}=\phi(m)$, $x(m_{i}^{2},m_{i}^{3},m_{j,j}^{1})\not=m_{i}^{3}$. By Claim 6, $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right)$. By Claim 6, we know that for any agent $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right),$ we have $x(m_{i}^{2},m_{i}^{3},\hat{m}_{j,j}^{1})\not=m$, for every $\hat{m}_{-i}$ such that $\sigma(m_{i},\hat{m}_{-i})>0$. In addition, $\hat{m}_{j,j}^{1}=\theta_{j}$ for every $j$ such that $j\not\in\mathcal{J}\left(m_{i}^{2},m_{i}^{3}\right).$ Hence, from agent $i$’s perspective, conditional on playing $m_{i},$ he is challenged with probability $1.$ Thus, agent $i$ is penalized $2\eta$ due to (22)). Consider a deviation to $\tilde{m}_{i}$ which is the same as $m_{i}$ except that $\tilde{m}_{i}^{2}=\theta$. Note that from agent $i$’s perspective, dictator lotteries are triggered with probability one, hence, all the agents report the truth in the first reports. Then, agent $i$ avoids paying the penalty $2\eta$ for being challenged, while the potential loss from allocation is bounded by $\eta,$ the loss from (17) is bounded by $\eta.$ Therefore, it is a profitable deviation. This contradicts to that $\sigma$ is an equilibrium. ∎
Claim 8.
For any pair of agents $i$ and $j$ such that $i\neq j$, message profile $m\insupp$$\left(\sigma\right)$ such that $m_{j}^{3}=\phi(m)$, if $x(m_{j}^{2},m_{j}^{3},m_{j,j}^{1})\neq m_{j}^{3}$, then we have $c_{i,j}>0$, $m_{i,j}^{1}=\theta_{j}$ and $m_{i,i}^{1}=\theta_{i}$.
Proof.
First, we fix an arbitrary correlated equilibrium $\sigma$. Given $m_{i}$ we collect all the opponents’ message profile together with $m_{i}$ agent $i$ knows that agent $j$ is challenged by himself:
$E_{j}(m_{i})=\{(m_{i},\tilde{m}_{-i}):g_{j,k}(m_{i},\tilde{m}_{-i})\neq\tilde{m}_{j}^{3}=\phi(m_{i},\tilde{m}_{-i})\text{ for some }k\text{ and }\sigma(m_{i},\tilde{m}_{-i})>0.\}$
From the condition in Claim 8, we know that $E_{j}(m_{i})\neq\varnothing$. By the proper scoring rule $sc_{i,j}$, $m_{i,j}^{4}=\sum_{\tilde{m}_{-i}\in E_{j}(m_{i})}\sigma_{-i}(\tilde{m}_{-i}|m_{i})$. Conditional on $m_{i}$, dictator lotteries are triggered with positive probability; hence by Lemmas 2.4 and 2.9, $m_{i,i}^{1}=\theta_{i}$. In addition, when agent $j$ is challenged by himself, we have $m_{j,j}^{1}=\theta_{j}$. Hence, according to $\tau_{i,j}^{2}$ (see (25)), we have $m_{i,j}^{1}=\theta_{j}.$
Claim 9.
For every agent $j$, message profile $m\insupp$$\left(\sigma\right)$ such that $m_{j}^{3}=\phi(m)$, we have $x(m_{j}^{2},m_{j}^{3},m_{k,k}^{1})=m_{j}^{3}$ for every agent $k$.
Proof.
By Claim 7, we know that for every $k\neq j$, Claim 9 holds. It remains to show that Claim 9 holds for $k=j$. Suppose there exists $m\insupp$$\left(\sigma\right)$, such that $m_{j}^{3}=\phi(m)$, if $x(m_{j}^{2},m_{j}^{3},m_{j,j}^{1})\neq m_{j}^{3}$. By Lemmas 2.4 and 2.9, $m_{j,j}^{1}=\theta_{j}$. Now we show that for agent $j$ who reports $m_{j}$ with $m_{j,j}^{2}\not=\theta_{j}$, it is strictly better for agent $j$ to deviate to $\tilde{m}_{j}$ which is identical to $m_{j}$ but $\tilde{m}_{j,j}^{2}=\theta_{j}$. Specifically, due to Claim 8, for every $\tilde{m}_{-j}$ such that $\sigma(m_{j},\tilde{m}_{-j})>0$, for every agent $i\neq j$, we have $\tilde{m}_{i,j}^{4}>0$, $\tilde{m}_{i,j}^{1}=\theta_{j}$. By $\tau^{3}$ in (28), we know that the deviation is profitable. Thus, it contradicts the hypothesis that $x(m_{j}^{2},m_{j}^{3},m_{j,j}^{1})\neq m_{j}^{3}$. ∎
Claim 10.
For every $m\in\text{supp}(\sigma)$, $g(m)=f(\theta)$, and $\tau_{i,j}(m)=0$ for every agent $i$ and $j$.
Proof.
By Claim 6 and Claim 9, we know that for every $m\in\text{supp}(\sigma)$, $g(m)\in F(\theta)$, and $\tau_{i,j}^{1}=0$ for every agent $i$ and $j$. Due to the construction of the proper scoring rule $sc$, we know that $m_{i,j}^{4}=0$ for every agent $i$ and $j$. Hence $\tau_{i,j}^{2}=\tau_{i,j}^{3}=0$ for every agent $i$ and $j$. Hence, we achieve implementation with no transfers incurred. ∎
2.8. Related Literature
(Aumann, 1974) first proposed the concept of correlated equilibrium (CE) as a generalization of the concept of independent randomization among agents and (Hart and Schmeidler, 1989) were the first to offer a direct proof of existence for CE. An important feature of correlated equilibria is that they can be derived as the result of several adaptive procedures. The first of these were proposed by (Foster and Vohra, 1997) in which agents forecast other’s play and play best responses to this forecast - this eventually yields a correlated equilibrium. (Fudenberg and Levine, 1999) show that a variant of fictitious play which satisfies a condition called conditional universal consistency also yields a CE. We build on the work of (Hart and Mas-Colell, 2000) who show (for finite N player games) that the regret matching heuristic yields CE of the underlying game. (Stoltz and Lugosi, 2007) generalize this result to infinite but convex and compact strategy sets. In principle our implementation result applies to each of these adaptive procedures and we select regret-matching for concreteness.
(Maskin, 1999) pioneered the concept of implementation of a social choice function (SCF) in Nash equilibria and derived the appropriate characterization, showing that a monotonicity condition was necessary and almost sufficient for implementability in Nash equilibria. Later results studied implementation in various solution concepts such as subgame perfect Nash equilibria (Moore and Repullo, 1988), iteratively undominated strategies (Abreu and Matsushima, 1992), and rationalizability (Bergemann et al., 2011). Classical mechanisms in this literature often use a device called an “integer game” which allows the designer to eliminate many undesirable message profiles by augmenting them with an integer so that the agent with the highest integer can dictatorially pick the outcome. This creates a “race to the top” and thus such games possess no equilibria. These “tail chasing devices” are critiqued in both (Jackson, 1992) and (Moore, 1992). Further, mechanisms which use integer games cannot be simulated (since they have an infinite message space) although a variant called a modulo game (with a finite message space) can be implemented in simulations. However, a modulo game yields undesirable mixed strategy equilibria, so that finite mechanisms are best suited for simulations. Recent papers have attempted to place the theory of implementation on a more theoretically sound footing by eschewing the use of integer games (Chen et al., 2022, 2021; Fehr et al., 2021), but general results for implementation in correlated equilibria have not yet been derived. This paper presents, to our knowledge, the first characterization of SCFs which can be implemented in CE and also presents a well behaved mechanism which does not use integer games.
Most recently, (Pei and Strulovici, 2025) show that the mechanism in (Chen et al., 2021) robustly implements SCFs satisfying maskin-monotonicity* by invoking Proposition 3.2 in (Kajii and Morris, 1997) which shows that a unique correlated equilibrium must be a robust equilibrium. They also present mechanisms which robustly implement SCFs which do not satisfy maskin monotonicity by relying on costly information. Relatedly, the mechanism proposed here implements SCFs satisfying maskin monotonicty (which is weaker than maskin monotonicity*) as the outcome of a unique correlated equilibrium of the game induced by the mechanism, and thus also robustly implements.
3. Simulation: Bilateral Trade
3.1. Environment
A seller $S$ has an object for sale to a buyer $B$. The seller can have a low or a high cost of production while the buyer may have a low or high valuation for the product. The costs for the seller are given by $c^{H}=8$ and $c^{L}=2$, while the valuations for the buyer are given by $v^{H}=20$ and $v^{L}=12$. Since each of the players in the game can have one of two types, there are four possible states of the world. The designer can impose transfers and hence the set of outcomes $A$ is the set of triplets $\left(q,t_{B},t_{S}\right)$ with $q\in\left[0,1\right]$ representing the amount of the good being traded, $t_{B}$ is the price paid by $B$ and $t_{S}$ is the payment received by $S$. For any outcome $\left(q,t_{B},t_{S}\right),$ $B$’s utility is $u_{B}=qv+t_{B}$ when the good quality is $v,$ and the seller’s utility is $u_{S}=t_{S}-qc$. The desired allocation (SCF) is shown in Table 1.
Table 1. Social Choice Function (SCF) Outcomes by State
State of the World SCF Outcome
Buyer Type Seller Type Quantity ($q$) Buyer Payment ($t_{B}$) Seller Payment ($t_{S}$)
L L 1 -6 6
H H 1 -10 10
H L 1 -10 10
L H 1 -10 10
3.2. Test Allocations
For the SCF to be implementable, it must satisfy Maskin Monotonicity. This condition ensures that if agents misreport the state of the world, at least one agent (a “whistle-blower”) will find it in their interest to challenge the lie.
Specifically, if the true state is $\theta$ but agents report a lie $\theta^{\prime}$, there must be a test allocation from a challenge scheme, denoted $x(\theta^{\prime},\theta_{i})$ as in Lemma 2.5, for a whistle-blowing agent $i$ (whose type differs between $\theta$ and $\theta^{\prime}$) such that two conditions are met:
  1. (1)
    No False Alarms: In the lie state $\theta^{\prime}$, the agent weakly prefers the original outcome to the test allocation. That is, $u_{i}(f(\theta^{\prime}),\theta^{\prime}_{i})\geq u_{i}(x(\theta^{\prime},\theta_{i}),\theta^{\prime}_{i})$.
  2. (2)
    Incentive to Expose Lie: In the true state $\theta$, the agent strictly prefers the test allocation to the outcome from the lie. That is, $u_{i}(x(\theta^{\prime},\theta_{i}),\theta_{i})>u_{i}(f(\theta^{\prime}),\theta_{i})$.
For the bilateral trade simulation, the challenge scheme $\{x(\theta^{\prime},\theta_{i})\}$ is defined with the following key test allocations:
  • When the Buyer is the whistle-blower (true type $v^{L}$, lie type $v^{H}$), the test allocation is $x(\theta^{\prime},v^{L})=(0.5,-2,2)$.
  • When the Seller is the whistle-blower (true type $c^{H}$, lie type $c^{L}$), the test allocation is $x(\theta^{\prime},c^{L})=(0.5,-3,3)$.
To demonstrate their validity, we will show that for the true state of the world being (L, H), a whistle-blower exists for every possible lie.
3.2.1. Lie State: (L, L)
If agents report $\theta^{\prime}=(L,L)$, the outcome is f(L,L)=(1,6,6)f(L,L)=(1,-6,6). The Seller’s true type ($c^{H}$) differs from the lie ($c^{L}$), making them the whistle-blower. We check the two conditions for the Seller using the test allocation $x((L,L),c^{H})=(0.5,-3,3)$.
  • Condition 1 (No False Alarms): We check if $u_{S}(f(L,L),c^{L})\geq u_{S}(x((L,L),c^{H}),c^{L})$.
    • $u_{S}(f(L,L),c^{L})=6-(2\times 1)=4$.
    • $u_{S}(x((L,L),c^{H}),c^{L})=3-(2\times 0.5)=2$.
    • The condition $4\geq 2$ holds.
  • Condition 2 (Incentive to Expose): We check if $u_{S}(x((L,L),c^{H}),c^{H})>u_{S}(f(L,L),c^{H})$.
    • $u_{S}(x((L,L),c^{H}),c^{H})=3-(8\times 0.5)=-1$.
    • $u_{S}(f(L,L),c^{H})=6-(8\times 1)=-2$.
    • The condition 1>2-1>-2 holds.
Both conditions are met, so the Seller can act as a whistle-blower.
3.2.2. Lie State: (H, H)
If agents report $\theta^{\prime}=(H,H)$, the outcome is f(H,H)=(1,10,10)f(H,H)=(1,-10,10). The Buyer’s true type ($v^{L}$) differs from the lie ($v^{H}$), making them the whistle-blower. We check the two conditions for the Buyer using the test allocation $x((H,H),v^{L})=(0.5,-2,2)$.
  • Condition 1 (No False Alarms): We check if $u_{B}(f(H,H),v^{H})\geq u_{B}(x((H,H),v^{L}),v^{H})$.
    • $u_{B}(f(H,H),v^{H})=(20\times 1)-10=10$.
    • $u_{B}(x((H,H),v^{L}),v^{H})=(20\times 0.5)-2=8$.
    • The condition $10\geq 8$ holds.
  • Condition 2 (Incentive to Expose): We check if $u_{B}(x((H,H),v^{L}),v^{L})>u_{B}(f(H,H),v^{L})$.
    • $u_{B}(x((H,H),v^{L}),v^{L})=(12\times 0.5)-2=4$.
    • $u_{B}(f(H,H),v^{L})=(12\times 1)-10=2$.
    • The condition $4>2$ holds.
Both conditions are met, so the Buyer can act as a whistle-blower.
3.2.3. Lie State: (H, L)
If agents report $\theta^{\prime}=(H,L)$, the outcome is f(H,L)=(1,10,10)f(H,L)=(1,-10,10). Here, both players’ true types differ from their reported types. We only need one to be a whistle-blower. Let’s check the Buyer (type change H $\rightarrow$ L) using the test allocation $x((H,L),v^{L})=(0.5,-2,2)$. The calculations are identical to the previous case, as the lie outcome is the same.
  • Condition 1 ($10\geq 8$) holds.
  • Condition 2 ($4>2$) holds.
The Buyer can act as a whistle-blower. Since a valid whistle-blower exists for every possible lie, the SCF is implementable for the true state of (L, H).
3.3. Dictator Lotteries
The mechanism also employs “dictator lotteries” to ensure that agents have a strict incentive to report their own type truthfully in certain parts of their message[cite: 50]. A set of dictator lotteries $\{y_{i}(\theta_{i})\}$ is valid if an agent $i$ with true type $\theta_{i}$ strictly prefers the lottery associated with their true type over the lottery for any other type $\theta^{\prime}_{i}$[cite: 51]. That is, for any $\theta_{i}\neq\theta^{\prime}_{i}$, it must be that $u_{i}(y_{i}(\theta_{i}),\theta_{i})>u_{i}(y_{i}(\theta^{\prime}_{i}),\theta_{i})$.
The lotteries used in the simulation are as follows:
3.3.1. Buyer’s Dictator Lotteries
The two dictator lotteries available for the buyer are:
  • If buyer’s reported type is H: $y_{B}(H)=(1,-15,15)$.
  • If buyer’s reported type is L: $y_{B}(L)=(0,0,0)$.
Validity Check:
  • When the Buyer’s true type is H ($v^{H}=20$):
    • Utility from true report $y_{B}(H)$: $u_{B}=(20\times 1)-15=5$.
    • Utility from false report $y_{B}(L)$: $u_{B}=(20\times 0)+0=0$.
    • The condition $5>0$ holds.
  • When the Buyer’s true type is L ($v^{L}=12$):
    • Utility from true report $y_{B}(L)$: $u_{B}=(12\times 0)+0=0$.
    • Utility from false report $y_{B}(H)$: $u_{B}=(12\times 1)-15=-3$.
    • The condition 0>30>-3 holds.
The lotteries are valid for the Buyer.
3.3.2. Seller’s Dictator Lotteries
The two dictator lotteries available for the seller are:
  • If seller’s reported type is H: $y_{S}(H)=(0,0,0)$.
  • If seller’s reported type is L: $y_{S}(L)=(1,-4,4)$.
Validity Check:
  • When the Seller’s true type is H ($c^{H}=8$):
    • Utility from true report $y_{S}(H)$: $u_{S}=0-(8\times 0)=0$.
    • Utility from false report $y_{S}(L)$: $u_{S}=4-(8\times 1)=-4$.
    • The condition 0>40>-4 holds.
  • When the Seller’s true type is L ($c^{L}=2$):
    • Utility from true report $y_{S}(L)$: $u_{S}=4-(2\times 1)=2$.
    • Utility from false report $y_{S}(H)$: $u_{S}=0-(2\times 0)=0$.
    • The condition $2>0$ holds.
The lotteries are also valid for the Seller.
3.4. Regret Minimization
The reader is referred to (Hart and Mas-Colell, 2000), (Hart and Mas-Colell, 2003), and (Hart, 2005) for a comprehensive description of the regret minimization heuristic and a theoretical proof of the claim that when players use the regret minimization heuristic to play a game, the long run distribution of the play eventually converges to a correlated equilibrium of the game. For completeness, we provide a very brief overview of the regret minimization heuristic here.
Consider a player $i$ with a set of messages $M_{i}$ in a repeated game setting over $T$ periods. Let $m_{i}^{t}\in M_{i}$ denote the message chosen by player $i$ in period $t$, and $m_{-i}^{t}$ denote the messages chosen by all other players in period $t$. Let $u_{i}(m_{i},m_{-i})$ be the payoff function of player ii. [^7] The reader is reminded that the utility function $u(m_{i},m_{-i})$ comprises the utility from the outcome, denoted by $v(g(m_{i},m_{-i}))$, and any transfers the mechanism might prescribe.
3.4.1. Regret Calculation
For each message $m_{i}^{\prime}\in M_{i}$, the regret of not having played $m_{i}^{\prime}$ up to period $T$ is defined as:
$R_{i}^{T}(m_{i}^{\prime})=\frac{1}{T}\sum_{t=1}^{T}\left[u_{i}(m_{i}^{\prime},m_{-i}^{t})-u_{i}(m_{i}^{t},m_{-i}^{t})\right],$
where $u_{i}(m_{i}^{\prime},m_{-i}^{t})$ is the hypothetical payoff player $i$ would have received had they played $m_{i}^{\prime}$ in period $t$, while holding $m_{-i}^{t}$ fixed.
3.4.2. Strategy Update Rule
Based on the calculated regrets $\{R_{i}^{T}(m_{i}^{\prime}):m_{i}^{\prime}\in M_{i}\}$, the player updates their probability distribution over messages. A common rule is to increase the probability of messages with positive regret. For example:
$p_{i}^{T+1}(m_{i}^{\prime})=\frac{\max\{R_{i}^{T}(m_{i}^{\prime}),0\}}{\sum_{m_{i}^{\prime\prime}\in M_{i}}\max\{R_{i}^{T}(m_{i}^{\prime\prime}),0\}},$
where $p_{i}^{T+1}(m_{i}^{\prime})$ denotes the probability of choosing message $m_{i}^{\prime}$ in the next period.
If all regrets are non-positive, the player may revert to a uniform or prior distribution over messages.
3.4.3. Regret Minimization Objective
The player’s objective is to minimize their average regret over time, defined as:
$\overline{R}_{i}^{T}=\max_{m_{i}^{\prime}\in M_{i}}R_{i}^{T}(m_{i}^{\prime}).$
A regret-minimizing algorithm ensures that $\overline{R}_{i}^{T}\rightarrow 0$ as $T\rightarrow\infty$, implying no significant regret for not having played any single message.
3.5. Simulation Details and Results
For ease of comparison across mechanisms, in the following simulations, the state of the world is set to $(L,H)$ and then two agents (the buyer and the seller) play the game induced by the mechanism detailed above using regret matching. The test allocations are chosen as detailed in Section 3.2. The initial strategies involve randomizing uniformly over the available messages. In about $200$ iterations, the dynamics are seen to have converged to the correlated equilibrium of the game, i.e. truthtelling, and no self-challenges. A sample simulation result is graphically represented in Figure 1.
As seen in Figure 1, the players gradually converge to playing the correlated equilibrium strategy with large probabilities despite only following the regret minimizing heuristic. Since the mechanism implements in correlated equilibria, this also yields the socially desirable outcome with a large probability. The panels in the bottom half of the figure show the evolution of the gains from trade with the growth of transfers induced by the mechanism, i.e. any transfers other than those required by the SCF itself.
3.6. Comparisons with other mechanisms
In (Hart and Mas-Colell, 2003), the authors note that “It is notoriously difficult to formulate sensible adaptive dynamics that guarantee convergence to Nash equilibrium”. (Chen et al., 2022) provides a mechanism which can be used to implement the SCF studied in this section in Nash equilibrium. We implement this mechanism (which we call the MAM mechanism) alongside regret dynamics using the same transfer scaling as the CE mechanism. The results are shown in Figure 2.
The mechanism takes longer (typically 2500 iterations or more) to converge to the correct actions, and incurs significantly more transfers in the process. Further, since it takes longer to find the optimal strategies, the social surplus is also significantly lower.
Allowing the dictator lotteries within the mechanism in (Chen et al., 2022) to always be on with probability $\epsilon$ provides an alternate mechanism in the spirit of (Abreu and Matsushima, 1992) which allows their mechanism to implement in correlated equilibria as well, although the implementation is virtual. We simulate this mechanism (which we call the AM92 mechanism) as well and present the results in Figure 3. This mechanism takes even longer to implement (of the order of 5000 iterations), and owing to the fact that the implementation is virtual, the dictator lotteries yield transfers with a small probability throughout the simulation, so that the transfers can be seen to dominate the social surplus. We conjecture that convergence is likely to be even slower with the actual mechanism in (Abreu and Matsushima, 1992) for a large number of rounds (K) since it would involve a larger message space. Further, the message space of the mechanism in (Abreu and Matsushima, 1992) depends on the number of rounds ($K$) which depends upon the degree of virtualness, $\epsilon$, whereas our mechanism remains the same regardless of the choice of $\epsilon$.
Figure 1. Simulation Results (CE Mechanism)
Figure 2. Simulation Results (MAM Mechanism)
Figure 3. Simulation Results (AM92 Mechanism)